OpenStudy (anonymous):

Find the product: (-7n + 3)(-9n +10)

OpenStudy (john_es):

You must do the product between all terms, like in this one, \[(a+b)(c+d)=a*b+a*d+b*c+b*d\]

OpenStudy (john_es):

Remember \[n*n=n^2\]